How do you graph y=2sin 1/3x?
i have to find the
Amplitude
period
phase shift
vertical shift
asymptotes
domain
and range

Answer by lwsshak3(11628)

How do you graph y=2sin 1/3x?
i have to find the
Amplitude
period
phase shift
vertical shift
asymptotes
domain
and range
**
Equation used to graph sin function: Asin(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B
For given sin function: y=2sin 1/3x
Amplitude=2
B=1/3
Period=2π/B=2π/(1/3)=6π
1/4 Period=3π/2
Phase-shift: none
Vertical shift: none
asymptotes: none
Domain: (-∞,∞) or all real numbers
Range: (-2,2) or between -2 and 2
..
Graphing for one period:
On the x-axis make tick marks at 3π/2, 3π, 9π/2 and 6π
Plot the following points: (0,0), (3π/2,2), (3π,0), (9π/2,-2) and (6π,0)
To graph more than one period, continue by just adding 1/4 period or 3π/2 to the x-coordinates.